OpenStudy (anonymous):

What is the value of 1/3x to the second power + 5.2y when x = 3 and y = 2? 12.4 13.4 14.4 16.4

OpenStudy (mathstudent55):

You mean this problem? Evaluate \(\dfrac{1}{3}x^2 + 5.2y \) for \(x = 3\) and \(y = 2\).

OpenStudy (mathstudent55):

The first step is to write the expression with x replaced by 3 and y replaced by 2. \(\dfrac{1}{3}x^2 + 5.2y =\) \( =\dfrac{1}{3}(3^2) + 5.2(2)\) Can you evaluate that expression? Remember PEMDAS, the order of operations.

OpenStudy (mathstudent55):

\(\dfrac{1}{3} \times 3^2 + 5.2 \times 2\) \( = \dfrac{1}{3} \times 9 + 10.4\) \(= 3 + 10.4\) \(= 13.4\)